OpenStudy (nathanjhw):

For a continuous function f(x), f(1) = 11, and integral (1,5) f'(x) =9. Determine f(5).

OpenStudy (nathanjhw):

2 9 11 18 20

OpenStudy (zarkon):

\[9=\int\limits_{1}^5f'(x)dx=f(5)-f(1)=f(5)-11\]

OpenStudy (nathanjhw):

So therefore f(5) = 20

OpenStudy (zarkon):

yes

OpenStudy (nathanjhw):

Thank you for the quick response!

OpenStudy (zarkon):

np